OpenStudy (anonymous):

Simplify the square root of -40+ the the square root of 90

OpenStudy (anonymous):

no real square root of -40

OpenStudy (anonymous):

are you working with reals or complex numbers?

OpenStudy (anonymous):

3sqrt(10)+i2sqrt(10)

OpenStudy (anonymous):

\[-\sqrt{40} + \sqrt{90}\]

OpenStudy (anonymous):

sqrt(10) is the answer then :)

OpenStudy (anonymous):

\[-2\sqrt{10}+3\sqrt{10}=\sqrt{10}\]